# A New Crossover Algorithm for LP Inspired by the Spiral Dynamic of PDHG

### v1

@@ 3.2 Practical Consideration
At the beginning of primal and dual push, we need to determine $B$ and $(D,N)$ to construct the auxiliary LP and OLS. For an IPM solution, which tends to stay away from bounds, the central path and strict complementarity are crucial for identifying non-basic variables. Identification is typically done by comparing the orders of magnitude of $\mathbf{x}$ and $\mathbf{c}-\mathbf{A}^{\top}\mathbf{y}$. In contrast, PDHG can benefit from the sparsity of its solution, induced by the projection, to more aggressively fix variables to their bounds.
@@
where $\bm{\delta}_{\mathbf{c}},\bm{\delta}_{\mathbf{b}}$ are independently and uniformly distributed over the interval $[0,1]$. The first normalized terms function as a center to stabilize the disturbance, preventing the unfixed non-basic variables from occasionally escaping boundaries.
@@
Currently, all the OLS problems are solved independently from scratch. Note that the coefficient matrices of two adjacent OLS problems differ in only a few columns or rows; thus, further development on reusing the factorization (probably via low-rank updates) may largely enhance the efficiency of our crossover to the level of the classic simplex-based crossover.
@@
IPMs and PDHG both have a strong ability to distinguish non-basic variables. IPMs can identify the strictly complementary solution pair with a sufficiently small duality gap (Andersen and Ye, 1996). PDHG can correctly classify variables as basic and non-basic in the scope of PDHG basis within finite iterations (Lu and Yang, 2024b).
@@ 4 Experiments
  • cuPDLP-C (Lu et al., 2023) can solve it to the relative tolerance of $81e-$ 8 within 600 seconds.
@@
We first solve the NETLIB LPs using cuPDLP-C inside COPT 7.1 (Ge et al., 2024) without GPU acceleration to the relative tolerance of $81e-$ 8. Then the optimal primal-dual solution is passed to the Julia implementation of our crossover algorithm. A crossover time limit of 300 seconds is set and the random seed is set to 0 for reproducibility.
@@
OLS solving to identify the spiral ray direction is a critical step in our crossover algorithm. For most instances, aside from six rare exceptions, the OLS is solved via an automatic selection between direct solver and LSMR based on the dimensions. Specifically, if the matrix $\mathbf{A}_{B}^{\top}$ or $\mathbf{A}_{D}$ in the OLS has fewer than 1000 rows or columns, we solve the OLS by direct factorization. Otherwise, we use the iterative solver LSMR with a relative tolerance of $161e-$ 16. The auxiliary LPs are again solved by cuPDLP-C.
For the six exceptions marked in blue, custom methods for OLS or auxiliary LPs are chosen. Due to occasional numerical issues, we always use direct solver or LSMR, which are labeled “d” and “*” respectively. The label “i” indicates the auxiliary LPs are solved by the IPM in COPT to avoid the timeout of cuPDLP-C.
@@ Results.
The numerical results are shown in Table LABEL:tbl:netlib, where “# supp. PDLP” is the number of support (i.e., components not on their bounds) of the initial solution, “# supp. COPT” and “# supp. cross” are the numbers of support of vertices from the crossover algorithm in COPT commercial solver and our crossover approaches. We also report the running time of our algorithm in the “time (sec)” column.
Our crossover overall succeeds in 93 out of 100 instances. The instance “cycle” only recovers a primal optimal vertex, but this vertex solution appears much sparser than the optimal vertex from COPT crossover. Here we still count it as a success, included in the 93 successful instances.
@@
Due to the instability of FOMs, the auxiliary LPs take a long time to solve for certain instances, sometimes even causing a timeout. The numerical difficulty with LPs and OLS is another factor contributing to the failure. However, despite failing to recover the optimal basis for these problems, the number of support of most primal solutions is still successfully reduced.
Table 1: Crossover results for 100 NETLIB instances. “a” means the basis is only primal optimal. “t” means crossover times out. “f” means crossover fails. “i” means the auxiliary LPs are solved by the IPM. For the OLS methods, “*” means always using LSMR, “d” means always using the direct solver, while other unmarked records automatically switch between direct solver and LSMR.
prob nRows nCols # supp. PDLP # supp. COPT # supp. cross time (sec)
25fv47 821 1571 600 583 584 11.63
80bau3b 2262 9799 1851 1758 1753 1.30
adlittle 56 97 61 45 44 0.01
afiro 27 32 14 13 14 0.01
agg 488 163 62 57 57 0.39
agg2 516 302 141 120 121 5.32
agg3 516 302 144 124 125 5.71
bandm 305 472 306 294 294 0.54
beaconfd 173 262 89 89 89 0.04
blend 74 83 56 54 54 0.01
bnl1 643 1175 689 451 448 11.33
bnl2 2324 3489 1519 1171 1168 40.05
boeing1 351 384 207 196 195 0.78
boeing2 166 143 69 55 55 0.04
bore3d 233 315 126 126 126 0.06
brandy 220 249 135 134 134 0.12
capri 271 353 246 220 225 0.26
cre-a 3516 4067 579 492 484 23.35d
cre-c 3068 3678 590 508 503 43.49d
cycle 1903 2857 994 224 105 36.27a
czprob 929 3523 924 866 866 27.00
d2q06c 2171 5167 1612 1543 1541 t
d6cube 415 6184 136 115 101 13.85
degen2 444 534 207 207 207 0.84
degen3 1503 1818 640 637 637 153.40
e226 223 282 127 127 127 0.17
etamacro 400 688 292 271 272 0.59
fffff800 524 854 357 312 315 2.59
finnis 497 614 259 227 229 1.61
fit1d 24 1026 12 12 12 0.01
fit1p 627 1677 634 627 627 7.79
fit2d 25 10500 22 20 20 0.01
fit2p 3000 13525 3004 2997 2997 5.47
forplan 161 421 83 83 83 1.69
ganges 1309 1681 1278 1175 1176 2.89i
gfrd-pnc 616 1092 349 336 336 1.32
greenbeb 2392 5405 1048 937 942 135.27
grow15 300 645 533 299 300 0.31
grow22 440 946 849 440 440 0.76
grow7 140 301 237 140 140 0.04
israel 174 142 80 70 69 0.63
kb2 43 41 27 27 27 0.01
ken-07 2426 3602 2236 2234 2234 1.44
lotfi 153 308 126 99 98 0.02
maros-r7 3136 9408 3136 3136 3136 0.39
maros 846 1443 345 340 270 5.58
modszk1 687 1620 666 666 666 10.68
nesm 662 2923 726 550 545 46.11i*
osa-07 1118 23949 357 355 355 0.20
osa-14 2337 52460 873 781 781 f
osa-30 4350 100024 1733 1536 1536 f
pds-02 2953 7535 1379 332 313 15.32
perold 625 1376 580 546 544 61.55
pilot.ja 940 1988 789 681 682 t
pilot 1441 3652 1299 1287 1298 t
pilot4 410 1000 374 364 362 104.69
pilot87 2030 4883 1856 1840 1839 t
pilotnov 975 2172 1809 708 680 f
qap12 3192 8856 2940 2462 2446 93.34
qap8 912 1632 656 466 418 9.54
recipe 91 180 24 24 24 0.01
sc105 105 103 85 85 85 0.02
sc205 205 203 184 184 184 0.09
sc50a 50 48 42 42 42 0.01
sc50b 50 48 48 48 48 0.07
scagr25 471 500 317 307 307 0.57
scagr7 129 140 98 97 97 0.01
scfxm1 330 457 248 231 231 0.15*
scfxm2 660 914 501 473 471 3.04
scfxm3 990 1371 754 711 711 12.42
scorpion 388 358 245 245 245 0.17
scrs8 490 1169 276 276 276 0.29
scsd1 77 760 31 7 11 0.02
scsd6 147 1350 182 63 56 0.03
scsd8 397 2750 551 147 143 0.21
sctap1 300 480 263 164 168 0.20*
sctap2 1090 1880 811 562 562 17.23
sctap3 1480 2480 1038 731 731 50.74
seba 515 1028 438 438 438 2.32
share1b 117 225 95 94 94 0.03
share2b 96 79 52 48 48 0.05
shell 536 1775 391 383 383 0.57
ship04l 402 2118 261 260 260 0.24
ship04s 402 1458 281 280 280 0.27
ship08l 778 4283 422 422 422 1.42
ship08s 778 2387 447 447 447 1.69
ship12l 1151 5427 707 706 706 4.67
ship12s 1151 2763 728 728 728 4.82
sierra 1227 2036 373 361 361 6.61
stair 356 467 350 349 349 0.76
standata 359 1075 71 50 50 0.03
standgub 361 1184 71 50 50 0.03
standmps 467 1075 190 174 174 0.26
stocfor1 117 111 69 69 69 0.01
stocfor2 2157 2031 1267 1267 1267 15.84
truss 1000 8806 802 691 689 17.45
tuff 333 587 165 122 113 0.76
vtp.base 198 203 55 55 55 0.03
wood1p 244 2594 39 39 39 1.49
woodw 1098 8405 696 550 550 3.31
@@ 5 Conclusions and Discussions
In this paper, we formally derive the spiral behavior of PDHG for solving LP. The spiral behavior can be decomposed by a rotation and a forward movement that are orthogonal to each other. This explicit formula and geometric perspective provide insights into the recent findings of the algorithm, including the infeasibility detection, the two-stage behavior, the local linear rate, etc.
Inspired by the spiral of PDHG, we propose a new randomized crossover approach that is distinct from the traditional crossover approach based on the simplex algorithm. Our crossover moves along the PDHG spiral ray, which can be calculated via solving OLS. Our numerical experiments demonstrate the effectiveness of the proposed approach.
Looking ahead, several future directions are worth exploring to enhance this crossover scheme. Firstly, we can reuse the factorization in the OLS solving, potentially speeding up the algorithm significantly. One potential enhancement is implementing a warm start for PDHG, which could speed up auxiliary LPs, although a theoretical guarantee for this remains to be established. Another promising direction involves developing a GPU-friendly OLS solver, which would allow the crossover method to leverage GPU parallelization. Additionally, designing a strategic perturbation could direct the solution toward a desired vertex more efficiently. If this perturbation is well-calibrated to avoid making the LP unbounded or infeasible, it could enable a purely PDHG-based, matrix-free crossover, making it particularly suitable for handling large-scale LPs.
Acknowledgement
This work was mostly done when the first author visited the second author at the University of Chicago, Booth School of Business. The authors thank the Booth School of Business for funding this visit. The authors also would like to thank Qi Huangfu for helpful discussions on the crossover algorithms in commercial LP solvers. The first author is partially supported by NSFC [Grant NSFC-72225009, 72394360, 72394365]. The second author is partially supported by AFOSR FA9550-24-1-0051.
@@ Appendix A Proof of Theorem 7
where $\bm{\Sigma}$ has $r=\text{rank}(\mathbf{A}_{B})$ non-zero diagonal elements, and $\mathbf{U},\mathbf{V}$ are two unitary matrices. The first $r$ columns of $\mathbf{U}$ and the remaining respectively form the basis of the image of $\mathbf{A}_{B}$ and the kernel of $\mathbf{A}_{B}^{\top}$. Analogously, the first $r$ columns of $\mathbf{V}$ and the remaining respectively form the basis of the image of $\mathbf{A}_{B}^{\top}$ and the kernel of $\mathbf{A}_{B}$.
@@
to obtain several $2\times 2$ blocks and contingent ones on the diagonal. Moreover, since $\mathbf{A}_{B}$ is a submatrix of $\mathbf{A}$ and $\eta\|\mathbf{A}\|_{2}<1$, we still have
@@
For one $2\times 2$ block $\mathbf{J}=\begin{bmatrix}1&\eta\sigma\\ -\eta\sigma&1-2\eta^{2}\sigma^{2}\end{bmatrix}$ with $\eta\sigma\in(0,1)$, it is easy to prove that $\lim_{k\to+\infty}\mathbf{J}^{k}=\mathbf{0}$ by verifying that the modulus of its eigenvalues is strictly less than 1. Therefore, we have
@@
where the last equation utilizes the property of Moore–Penrose inverse that $\mathbf{A}_{B}=\mathbf{A}_{B}\mathbf{A}_{B}^{\dagger}\mathbf{A}_{B}$.

### v2

@@ 3.2 Practical Consideration
Auxiliary LP heuristic.
To reduce the required number of OLS subproblems, inspired by Ge et al. (2025b), we adopt auxiliary LPs as heuristic in primal and dual push to sparsify the original solution. In primal push, we build an auxiliary LP with random cost perturbation $\tilde{\mathbf{c}}_{B}$,
$\min$ $\tilde{\mathbf{c}}_{B}^{\top}\mathbf{x}_{B}$ (17)
$\operatornamewithlimits{s.t.}$ $\mathbf{A}_{B}\mathbf{x}_{B}=\mathbf{b}$
$\mathbf{x}_{B}\geq\mathbf{0},$
which can be solved with PDLP or IPMs. This auxiliary LP has multiple optimal solutions w.p. 0 (see Lemma 1 in Ge et al. (2025b)), so with proper perturbation to avoid the unbounded case, we have the unique optimal solution $\mathbf{x}_{B}$ to (17). Then $(\mathbf{x}_{B},\mathbf{0}_{E})$ can be the primal part of an optimal vertex of (1).
For the auxiliary LP in dual push, we sum up the non-negative dual slacks as a penalty to activate more dual constraints.
$\min$ $\mathbf{1}^{\top}(\mathbf{c}_{N}-\mathbf{A}_{N}^{\top}\mathbf{y})$ (18)
$\mathbf{A}_{N}^{\top}\mathbf{y}\leq\mathbf{c}_{N}.$
Due to the numerical residuals in practice, we may not correctly fix all the variables on bounds at once. To reduce the negative impact of the issue, the amplitude of disturbance should be controlled, see Mehrotra (1991); Ge et al. (2025b) for more specific discussions.
@@
[…]
@@
where $\bm{\delta}_{\mathbf{c}},\bm{\delta}_{\mathbf{b}}$ are independently and uniformly distributed over the interval [0,1][0,1][ 0 , 1 ]. The first normalized terms function as a center to stabilize the disturbance, preventing the unfixed non-basic variables from occasionally escaping boundaries.
@@
Note that to obtain the ray directions of PDHG (7), if the direct QR factorization is applied in the OLS subproblems (15) and (16), we need to factorize $\mathbf{A}_{B}^{\top}$ and $\mathbf{A}_{D}$ respectively. An alternative practice is to utilize the property of Moore–Penrose inverse that $(\mathbf{A}_{B}^{\top}\mathbf{A}_{B})^{\dagger}\mathbf{A}_{B}^{\top}=\mathbf{A}_{B}^{\dagger}=\mathbf{A}_{B}^{\top}(\mathbf{A}_{B}\mathbf{A}_{B}^{\top})^{\dagger}$. We can solve
$\min\ \frac{1}{2}\|\mathbf{A}_{B}\mathbf{d}_{B}-\mathbf{A}_{B}\tilde{\mathbf{c}}_{B}\|_{2}^{2},$ (21)
and calculate the primal direction $(\bm{\delta}_{\mathbf{x}})_{B}$ as
$(\bm{\delta}_{\mathbf{x}})_{B}=-\eta\left[\mathbf{c}_{B}-\mathbf{A}_{B}^{\top}(\mathbf{A}_{B}\mathbf{A}_{B}^{\top})^{\dagger}\mathbf{A}_{B}\tilde{\mathbf{c}}_{B}\right]=-\eta\left(\mathbf{c}_{B}-\mathbf{d}_{B}\right).$
Consequently, we can focus solely on the factorization of $\mathbf{A}_{B}$.
Currently, all the OLS problems are solved independently from scratch. Note that the coefficient matrices of two adjacent OLS problems differ in only a few columns or rows; thus, further development on reusing the factorization (for example, via Givens rotations) may largely enhance the efficiency of our crossover to the level of the classic simplex-based crossover.
@@
IPMs and PDHG both have a strong ability to distinguish non-basic variables. IPMs can identify the strictly complementary solution pair with a sufficiently small duality gap (Andersen and Ye, 1996). PDHG can correctly classify variables as basic and non-basic in the scope of PDHG basis within finite iterations (Lu and Yang, 2024a).
@@ 4 Experiments
  • cuPDLP-C (Lu et al., 2023) can solve it to the relative tolerance of $1e-8$ within 600 seconds.
@@
[…]
@@
OLS solving to identify the spiral ray direction is a critical step in our crossover algorithm. In our experiments, all OLS subproblems are solved by direct QR factorization.
Auxiliary LP solving.
The auxiliary LPs in our crossover experiment are solved to the relative tolerance of $1e-8$ by cuPDLP-C without GPU acceleration. The heuristic is disabled by default.
Vertex solutions from crossover.
From Theorem 11, our crossover will provide us with a primal-dual optimal vertex solution w.p. 1 if the original solution is primal-dual optimal. In practice, due to the numerical residual of the original solution, the vertex from the crossover may sometimes be only primal (or dual) optimal but dual (or primal) infeasible, or even near-optimal but slightly primal-dual infeasible. Since simplex methods can be applied to efficiently eliminate this infeasibility after crossover, such vertices are acceptable.
@@ Results.
The complete numerical results of the 100 NETLIB instances are shown in Appendix F. For the original $1e-8$ solutions, our crossover returns 95 primal-dual optimal vertices, 1 primal optimal vertex, and 4 dual optimal vertices. For the original $1e-6$ solutions, our crossover returns 76 primal-dual optimal vertices, 9 primal optimal vertices, 10 dual optimal vertices, and 5 near-optimal vertices. The inaccuracy of the original solution results in a loss of optimality in the vertex solution produced by the crossover procedure.
@@
With the auxiliary LP heuristic for original $1e-8$ solutions, our crossover can successfully provide optimal or near-optimal vertex solutions in 93 instances. Compared to the results without the auxiliary LP heuristic, the number of OLS subproblems is significantly reduced, but extra LP solving time is introduced, resulting in a trade-off. Due to the instability of FOMs, the auxiliary LPs take a long time to solve for certain instances, sometimes even causing a timeout. However, despite failing to recover the optimal basis for these problems, the number of support of most primal solutions is still successfully reduced.
@@ 5 Conclusions and Discussions
In this paper, we formally derive the spiral behavior of PDHG for solving LP. The spiral behavior can be decomposed into a rotation and a forward movement that are orthogonal to each other. This explicit formula and geometric perspective provide insights into the recent findings of the algorithm, including the infeasibility detection, the two-stage behavior, the local linear rate, etc.
Inspired by the spiral of PDHG, we propose a new randomized crossover approach that is distinct from the traditional crossover approach based on the simplex algorithm. Our crossover moves along the PDHG spiral ray, which can be calculated via solving OLS subproblems. Our numerical experiments demonstrate the effectiveness of the proposed approach.
Looking ahead, several future directions are worth exploring to enhance this crossover scheme. Firstly, we can reuse the factorization in the OLS solving, potentially speeding up the algorithm significantly. One potential enhancement is implementing a warm start for PDHG, which could speed up auxiliary LPs, although a theoretical guarantee for this remains to be established. Another promising direction involves developing a GPU-friendly OLS solver, which would allow the crossover method to leverage GPU parallelization. Additionally, designing a strategic perturbation could direct the solution toward a desired vertex more efficiently. If this perturbation is well-calibrated to avoid making the LP unbounded or infeasible, the auxiliary LP heuristic could enable a purely PDHG-based, matrix-free crossover, making it particularly suitable for handling large-scale LPs.
Acknowledgments
This work was started when the first author visited the second author at the University of Chicago, Booth School of Business. The authors also would like to thank Qi Huangfu for helpful discussions on the crossover algorithms in commercial LP solvers. The first author is partially supported by NSFC [Grant NSFC-72225009, 72394360, 72394365]. The second author is partially supported by AFOSR FA9550-24-1-0051.
@@ Appendix A Proof of Theorem 7
where $\bm{\Sigma}$ has $r=\text{rank}(\mathbf{A}_{B})$ nonzero diagonal elements, and $\mathbf{U},\mathbf{V}$ are two unitary matrices. The first $r$ columns of $\mathbf{U}$ and the remaining respectively form the basis of the image of $\mathbf{A}_{B}$ and the kernel of $\mathbf{A}_{B}^{\top}$. Analogously, the first $r$ columns of $\mathbf{V}$ and the remaining respectively form the basis of the image of $\mathbf{A}_{B}^{\top}$ and the kernel of $\mathbf{A}_{B}$.
@@
to obtain several $\times 2$ 2 blocks and contingent ones on the diagonal. Moreover, since $\mathbf{A}_{B}$ is a submatrix of $\mathbf{A}$ and $\eta\|\mathbf{A}\|_{2}<1$, we still have
@@
For one $\times 2$ 2 block $\mathbf{J}=\begin{bmatrix}1&\eta\sigma\\ -\eta\sigma&1-2\eta^{2}\sigma^{2}\end{bmatrix}$ with $\eta\sigma\in(0,1)$, it is easy to prove that $\lim_{k\to+\infty}\mathbf{J}^{k}=\mathbf{0}$ by verifying that the modulus of its eigenvalues is strictly less than 1. Therefore, we have
@@
$=\mathbf{A}_{B}^{\top}(\mathbf{A}_{B}\mathbf{A}_{B}^{\top})^{\dagger}\mathbf{b}+\operatorname{Proj}_{\mathbf{A}_{B}\mathbf{x}=\mathbf{0}}(\mathbf{x}_{B}^{(0)}).$
@@
where the last equation utilizes the property of Moore–Penrose inverse that $\mathbf{A}_{B}=\mathbf{A}_{B}\mathbf{A}_{B}^{\dagger}\mathbf{A}_{B}$. ∎